OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9

OpenStudy (anonymous):

@ash2326 @hartnn

OpenStudy (ash2326):

\[4x-y=8\] you need to find y in terms of x, Subtract both sides by 4x

OpenStudy (anonymous):

y=8?

OpenStudy (ash2326):

\[4x-y=8\] Subtract 4x from both the sides \[4x-y-4x=8-4x\] \[-y=8-4x\] Now multiply the whole equation by -1

OpenStudy (anonymous):

y = -8 - (-4x)

OpenStudy (ash2326):

\[y=4x-8\] Correct, so you'll substitute this expression in second to solve